Calculate the Taylor Polynomial Tg for the function f(x) =In(x + 2) centered at a = -1 TT T Arial 3 (12pt) T
Answer #1

In Oct 2) = ( کے feuse I trust at Xtz نر foya (142) feug = رہ Sofaen او را - صور ره = (1-) = = = = = ( نمر = ( ( درا = - ScanTs (3) = 400+6413fc-1) + (8 +0% f (-1) t SCHUS 22 FC) + Ct !! LI (Oct 19 w f (-) f ( ( ) ۔ 15 2 Tr (x) = ott) (1) + (HD + Oxca e HamScanner To (X) = (CH) - 18 +12 + + 6+1) + 8 + 1)² + + OC tus CS